# HAIFIT: Human-Centered AI for Fashion Image Translation

### v1

@@ I Introduction
In the earlier research, methods like Sketch2Photo [2] synthesized realistic images by amalgamating object and background images derived from provided sketches. They laid the groundwork for deep learning techniques. The emergence of Generative Adversarial Networks (GAN) [3] marked a significant shift in this domain, with methods such as BicycleGAN [25], pix2pix [7], and others [6, 11, 9, 21, 22] gaining prominence. While these methods have demonstrated commendable results, preserving sketch details remains a challenge as shown in Fig. 1. This limitation significantly hampers designers’ ability to preview and enhance finer details and refinements within sketches. Recent years have witnessed the rise of diffusion models [16, 10, 23], however, these methods heavily rely on extensive data for learning and necessitate substantial computational resources. Consequently, they pose challenges in terms of training, inference, and fine-tuning, incurring high time and resource costs, which may not be user-friendly for designers.
@@
Figure 1: The performance of different methods [19, 10] in handling sketch details (zooming in for better visual).
@@
To address these challenging problems, we introduce HAIFIT (Human-Centered AI for Fashion Image Translation), a GAN-based creative generation network designed for sketch-to-image translation in fashion design. HAIFIT aims to empower designers to foster inspiration by preserving as much detail as possible from their sketches in the generated images. Specifically, HAIFIT focuses on converting clothing sketches into high-quality realistic clothing images using a pyramidal generative adversarial network. To facilitate this conversion, we have developed a customized Multi-scale Feature Fusion Encoder (MFFE) that maps the reference hand-drawn sketches and generated images into a shared latent space. MFFE comprises a Shallow Convolution Module (SCM) and an Abstract Feature Representation Module (AFRM), which enables the capture of long-term dependencies between different regions and improves the global correlation in feature representations. Furthermore, we enhance the fine-grained features among generated images of various scales by integrating the Cross-level Skip Connection (CSC) within the feature space across different layers. By employing multi-scale discriminators, we ensure the preservation of global structural integrity while emphasizing fine details, resulting in more realistic generated images. The key contributions of our work are summarised as follows:
@@
  • We propose an innovative sketch-to-image generation network for clothing fashion design. Our model leverages an LSTM-based MFFE to enhance the representation of global features, ensuring coherence and consistency of features in the latent space. Furthermore, our CSC strategically directs the model’s attention towards finer sketch details, contributing to an improved and more nuanced sketch feature representation.
  • We construct a clothing dataset from fashion designers, which includes most types of clothing sketch-image pairs. To the best of our knowledge, it is the first professional clothing sketch-image dataset.
@@
  • Our method demonstrates superior performance in extensive quantitative and qualitative experiments compared to various baselines.
Figure 2: The illustration of our HAIFIT architecture. Given $M$ paired sketch-image samples $\{(x_{i},y_{i})\in(X,Y)\}^{M}_{i=1}$, our goal is to maps images from hand-drawn sketches domain $X$ to realistic images domain $Y$.
@@ II-A Sketch-to-Image Generation
The task of sketch-to-image generation falls within the realm of image translation, representing a pivotal objective in both computer graphics and computer-human interaction. Isola et al. [7] pioneered a unified framework for image-to-image translation models, adaptable across various scenarios. For instance, it allows segmented images to be mapped to natural images or transforms edge maps into colored images, among other applications. Pix2piHD [19] utilized multi-scale generators and discriminators to achieve high-resolution image translation on this basis. Other methods [6, 1, 11, 9] integrate attention networks, novel adversarial loss functions, or distinct image features to improve image generation performance. Nevertheless, these approaches still encounter challenges in managing intricate strokes and rectifying synthetic structures to produce realistic images.
@@ II-B Generative Adversarial Networks
Generative Adversarial Network (GAN) was initially proposed by Goodfellow et al. [3], introducing a framework where the Generator and Discriminator engage in adversarial training to produce realistic data. Mirza et al. [13] expanded the GAN model to enable the generation of images with specific conditions, empowering users to control attributes like image categories. Radford et al. [14] extended GANs by employing convolutional neural networks, enhancing their ability to capture spatial image characteristics. Subsequent research [8] further broadens its capabilities, focusing on the stable generation of high-quality and diverse images. The continual advancement and evolution of GANs offer immense potential in the field of image generation. Compared to the recent diffusion models [16, 10, 23], GANs possess the advantages of low hardware costs and fast synthesis speed, thus continuing to play a key role in the design field.
@@ III Methodology
In this study, we present HAIFIT, as depicted in Fig. 2, which comprises two closely related components: the Multi-scale Feature Fusion Encoder (MFFE), and the multi-scale pyramid generator with Cross-level Skip Connection (CSC). The MFFE is used to get global contour features and abstract intent features. Then the feature maps are resized and concatenated at multiple scales before being fed into the multi-scale pyramid generator, which can produce realistic images with complete structures and exquisite details.
III-A Multi-scale Feature Fusion Encoder
@@
The Multi-scale Feature Fusion Encoder (MFFE) consists of two modules: the Shallow Convolutional Module (SCM) and the Abstract Feature Representation Module (AFRM). As illustrated in Fig. 3, the SCM employs a convolutional network to learn essential global contour information, such as the structure of the object. Simultaneously, inspired by the LSTM architecture  [5], we design the AFRM to learn long-range dependencies in key areas of the input sketch from various perspectives, which can effectively capture style features and geometric features of strokes. By combining these features, we can effectively capture objects’ complete structure and intricate textures within the latent space.
@@
As shown in Fig. 2, in the bottom layer, only the sketch features are inputted. In the other layers, we perform element-wise addition between the generated image features and the sketch features, and then input the resulting summed features into the MFFE. In the SCM we employ a sequence of convolutional layers, denoted as $Con$, to extract the global contour features and shape information $\widetilde{x}$ from the input sketch $x_{s}$. Here, $x_{s}\in\mathcal{R}^{b\times 3\times m\times m}$, $\widetilde{x}\in\mathcal{R}^{b\times 256\times m/4\times m/4}$, $b$ is the batch size, and $m$ depends on the network layer, so we can express this process as $\widetilde{x}=Con(x_{s})$. Specifically, the SCM extracts features in a blockwise manner, which will disrupt the continuity of strokes and result in pixel-level contour features. Since sketches are continuous and composed of multiple lines in different directions, it is crucial to consider the correlation between adjacent lines in order to effectively learn the line features of complex sketches. To address this issue, we utilize the AFRM to learn the feature $\widetilde{x}$. The AFRM consists of a two-layer LSTM network, capturing abstract drawing intent features from four directions: left, right, top, and bottom. After downsampling $\widetilde{x}$, we obtain the feature $\widetilde{x^{\prime}}\in\mathcal{R}^{b\times 256\times 4\times 4}$. We then swap feature positions after merging its second and third dimension features to obtain $\mathcal{R}^{4\times b\times 1\times 1024}$ and divide it into four parts, namely $\widetilde{x}_{left}$, $\widetilde{x}_{right}$, $\widetilde{x}_{top}$, and $\widetilde{x}_{bottom}$, with dimensions $\mathcal{R}^{b\times 1\times 1024}$. These features are then individually sequence inputted into the LSTM network, generating features $\overline{x}_{left}$, $\overline{x}_{right}$, $\overline{x}_{top}$, and $\overline{x}_{bottom}$ by recursively processing the input features from different directions. Next, we concatenate the features $\overline{x}_{left}$, $\overline{x}_{right}$, $\overline{x}_{top}$, and $\overline{x}_{bottom}$ to obtain the feature $\overline{x^{\prime}}\in\mathcal{R}^{4\times b\times 1\times 1024}$. Afterward, we restore the dimension of $\overline{x^{\prime}}$ to the dimension of $\widetilde{x^{\prime}}$. Finally, we apply a 1$\times$1 convolutional layer along the channel dimension to obtain the output feature $\overline{x}\in\mathcal{R}^{b\times 256\times 4\times 4}$.
@@
$\begin{split}\overline{x}_{dir}^{\prime}&=Cat(LSTM[\widetilde{x}_{left},\widetilde{x}_{right},\widetilde{x}_{top},\widetilde{x}_{bottom}])\\ \overline{x}_{rev}^{\prime}&=Cat(LSTM[\widetilde{x}_{bottom},\widetilde{x}_{top},\widetilde{x}_{right},\widetilde{x}_{left}])\\ \overline{x}&=Conv_{1\times 1}(Cat(\overline{x}_{dir}^{\prime},\overline{x}_{rev}^{\prime}))\end{split}$ (1)
@@
III-B A Pyramid Generator with Cross-level Skip Connection
@@
In order to effectively capture intricate details in the key areas of the sketch, we employ a pyramid generative learning approach with a Cross-level Skip Connection (CSC). Our model comprises a pyramid of generators $G_{n}\in\{G_{0},G_{1},\dots,G_{N}\}$, which are trained using a set of multi-scale images $x_{i}$ represented as ${x^{0}_{i},x^{1}_{i},\dots,x^{N}_{i}}$. Each generator $G_{n}$ is responsible for generating a realistic image $\overline{y}^{n}_{i}$. By adversarial training, the generator $G_{n}$ aims to deceive the associated discriminator $D_{n}\in\{D_{0},D_{1},\dots,D_{N}\}$, and the $D_{n}$ is trained to distinguish between the generated realistic images and real images. The generation of a realistic image sample starts at the coarsest scale $64\times 64$ and sequentially passes through all generators, progressing toward the finest scale $256\times 256$. At the coarsest scale, the MFFE takes a minibatch of the sketch feature maps $X_{n}$ as input. The effective receptive field at this level is typically half the height of the finer image. Each of the generators $G_{n}$ at finer scales adds details that are not generated by the previous scales. Therefore, in addition to the feature maps $X_{n}$, each multi-scale feature fusion encoder accepts an upsampled version of the generated image features from the coarser scale. Here, $X_{n}\in\mathcal{R}^{b\times 3\times m\times m}$, $Y_{n-1}\in\mathcal{R}^{b\times 3\times m/2\times m/2}$. This pyramid structure allows us to capture features at multiple scales and better handle the complex details of the sketch.
[…]
Figure 4: An overview of the Cross-level Skip Connection (CSC).
@@
The generator $G_{n}$ consists of a residual block and multiple convolutional layers. In the coarsest layer, the residual block comprises four convolutional layers. In addition, in the pyramid network, as the number of pixels increases from bottom to top, the residual block progressively adds one convolutional layer at each level. This progressive increment facilitates the learning of sketch feature representations. Through numerous experiments, it has been observed that incorporating the CSC between different layers in the pyramid network can enhance the quality of the generated images, which is illustrated in Fig. 4. In detail, We extract the output feature ${X^{\prime}}_{n-1}$ from the last layer of the residual block in generator $G_{n-1}$ and concatenate it with the output feature ${X}_{n}$ from the $n$-th layer of the MFFE. Subsequently, we use $\overline{X}_{n}$ as the input for generator $G_{n}$, where $\overline{X}_{n}\in\mathcal{R}^{b\times 256\times m\times m}$. In principle, the multi-level features of the same synthetic image should exhibit similarity. By employing cross-level feature operations, we can extract similar details from the features generated by different level generators and incorporate them as attention maps into the new feature map $X_{n}$. This facilitates image translation, resulting in more detailed and high-quality images.
@@
III-C Loss Function

### v2

@@ I Introduction
In the earlier research, methods like Sketch2Photo [4] synthesized realistic images by amalgamating object and background images derived from provided sketches. They laid the groundwork for deep learning techniques. The emergence of Generative Adversarial Networks (GAN) [6] marked a significant shift in this domain, with methods such as BicycleGAN [35], pix2pix [12], and others [11, 18, 15, 31] gaining prominence. Although these methods have showcased notable results, preserving sketch details remains a challenge, as illustrated in Fig. 1. Through our comparison of two classic algorithms based on GAN [29] and Diffusion Model (DM) [17], we discovered significant hurdles in retaining sketch details. Specific elements such as clothing necklines, pockets, sleeves, and other intricate details prove challenging to effectively express in the lines of the designer’s sketches. This limitation significantly hampers designers’ ability to preview and enhance finer details and refinements within sketches.
@@
Figure 1: The performance of different methodological approaches [29, 17] compared to ours in handling sketch details (zooming in for better visual).
In recent years, with the introduction of the diffusion model [9], a novel and improved image generation algorithm has swiftly emerged. Some works, such as DDIM [26], employ skip sampling to expedite the model’s inference process, while LDM [22] utilizes feature mapping to execute the diffusion process in the latent space, effectively mitigating the issue of high resource requirements for model training. Additionally, Palette [24] and BBDM [17] showcase the efficacy of diffusion models in image translation tasks. However, these methods heavily rely on extensive data for learning and demand substantial computing resources. Consequently, they still face significant challenges in model training, inference, and fine-tuning, resulting in high time and resource costs, which may not be user-friendly for designers.
@@
To address these challenges, we introduce HAIFIT (Human-Centered AI for Fashion Image Translation), a GAN-based creative generation network designed for sketch-to-image translation in fashion design. HAIFIT aims to empower designers to foster inspiration by preserving as much detail as possible from their sketches in the generated images. Specifically, HAIFIT focuses on converting clothing sketches into high-quality realistic clothing images using a pyramidal generative adversarial network. To facilitate this conversion, we have developed a customized Multi-scale Feature Fusion Encoder (MFFE) that maps the reference hand-drawn sketches and generated images into a shared latent space. MFFE comprises a Shallow Convolution Module (SCM) and an Abstract Feature Representation Module (AFRM), which enables the capture of long-term dependencies between different regions and improves the global correlation in feature representations. Furthermore, we enhance the fine-grained features among generated images of various scales by integrating the Cross-level Skip Connection Module (CSCM) within the feature space across different layers. By employing multi-scale discriminators, we ensure the preservation of global structural integrity while emphasizing fine details, resulting in more realistic generated images. The key contributions of our work are summarised as follows:
  • Methodology. To the best of our knowledge, we are the pioneers in treating sketch features as sequence inputs, ensuring consistency by inputting features from both forward and reverse feature sequences. This innovative approach enhances the representation of details in the generated images.
@@
  • Structure. We introduce the LSTM-based MFFE to enhance the representation of global features and ensure coherence and consistency in the latent space. Additionally, our CSCM strategically directs the model’s attention to finer sketch details, contributing to improved and more nuanced sketch feature representation.
  • Datasets. We construct a background-free dataset comprising fashion clothes and their corresponding sketches for cross-modal image translation tasks, filling the gap in this field.
@@
  • Performance. Our method demonstrates superior performance in extensive quantitative and qualitative experiments compared to various baselines.
@@ II-A Sketch-to-Image Generation
The task of sketch-to-image generation falls within the realm of image translation, a pivotal objective in computer graphics and computer-human interaction. In the domain of fashion clothing design, having an effective algorithm that can accurately generate diverse clothing images based on designers’ sketches is invaluable. Such algorithms enable designers to express their creative ideas effectively, save development time, and enhance design efficiency. Early methods, like Photosketcher [5], utilized local features for image retrieval and employed graphics techniques to segment and stitch images together to achieve sketch-to-image generation. Subsequently, with the development of convolutional neural networks [16, 27, 25, 7], deep learning-based methods have become the mainstream approach for achieving sketch-to-image generation. Isola et al. [12] introduced a unified framework for image-to-image translation models adaptable to various scenarios. This framework facilitates mapping segmented images to natural images or converting edge maps to color images, among other applications. Building upon this, Pix2piHD [29] utilized multi-scale generators and discriminators to achieve high-resolution image translation. Several other methods [11, 1, 18, 15] have improved image generation performance by incorporating attention networks, innovative adversarial loss functions, or different image features. However, these methods still encounter challenges in handling complex strokes and refining synthetic structures to produce realistic images. In summary, synthesizing an image that faithfully represents the content of a sketch while maintaining consistent styles is paramount.
@@ II-B Generative Adversarial Networks
The Generative Adversarial Network (GAN) comprises two components: the generator and the discriminator. It leverages adversarial training to enhance the generator’s ability to generate realistic images, yielding significant achievements in image generation, restoration, and various other domains. Initially proposed by Goodfellow et al. [6]. For generating realistic data, GANs have since undergone several extensions and improvements. Mirza et al. [20] extended the GAN to enable the generation of images with specific conditions, allowing users to control attributes such as image categories. Radford et al. [21] introduced convolutional neural networks into both the generator and discriminator, thereby enhancing GAN’s ability to capture spatial image features. Zhu et al. [34] achieved cross-domain image conversion by introducing a cycle consistency loss function, enabling realistic transformations from one domain to another, such as from horse to zebra. Karras et al. [14] achieved stable control over the style and diversity of generated images by introducing style variables. Brock et al. [3] further scaled up the GAN and improved training methods to generate high-resolution images. Additionally, methods like WGAN [2], PGGAN [13], and SAGAN [32] have introduced new optimization losses, multi-layer resolution training strategies, and self-attention mechanisms to achieve more stable GAN training and further enhance the quality and diversity of generated images. In general, the continuous development and evolution of GANs have unlocked significant potential in the field of image generation, transforming the generation of realistic images from a concept into reality. Compared to the recent diffusion models [24, 17], GANs possess the advantages of low hardware costs and fast synthesis speed, thus continuing to play a key role in the design field.
@@ III Methodology
Algorithm 1 The PGGAN Training Process.
0:  $n$: Number of multi-layers
0:  $G=\{G_{1},G_{2},...,G_{n}\}$: Generator networks
0:  $D=\{D_{1},D_{2},...,D_{n}\}$: Discriminator networks
0:  $I=\{I_{1},I_{2},...,I_{n}\}$: Image resolutions
0:  $k$: Number of training iterations
1:  $G\leftarrow G_{1}$, $D\leftarrow D_{1}$, $I\leftarrow I_{1}$
2:  Initialize $G_{1}$ and $D_{1}$ with random weights
3:  for $i=1$ to $n$ do
4:     Training $D_{i}$ for $k$ iterations with fixed $G_{i}$
5:     Training $G_{i}$ for $k$ iterations with fixed $D_{i}$
6:     if $I_{i}<I_{n}$ then
[…]
8:        Initialize new layers $G_{i+1}$ and $D_{i+1}$ for higher resolution
9:        Update weights of existing layers in $G$ and $D$
10:     end if
11:  end for
@@
In this study, we present HAIFIT, as depicted in Fig. 2, which is based on PGGAN [13] and comprises two core components: the Multi-scale Feature Fusion Encoder (MFFE), and the Cross-level Skip Connection Module (CSCM). The MFFE is used to get global contour features and abstract intent features. Then the feature maps are resized and concatenated at multiple scales before being fed into the multi-scale pyramid generator, which can produce realistic images with complete structures and exquisite details.
III-A Preliminaries
The traditional GANs [6] enhance the model’s generation capabilities by adversarially training the generator and discriminator. However, they still face significant challenges in terms of image generation quality and training process stability. To address these issues, PGGAN [13] introduces a progressive growth training method. This approach starts from low-resolution images and gradually increases the number of layers and image resolution of both the generator and discriminator. This progressive growth training process not only simplifies the model’s convergence to generate higher-quality and more flexible-resolution images, but also mitigates issues such as mode collapse and gradient disappearance during the training process. Consequently, we utilize the PGGAN as the framework for our approach.
The specific algorithmic process is outlined in Algorithm 1. We commence training the initial generator $G_{1}$ and initial discriminator $D_{1}$ from the initial resolution $I_{1}$, gradually increasing the resolution and continuing training by adding new $G_{i+1}$ and $D_{i+1}$ layers. This process continues until the final resolution $G_{n}$ and $D_{n}$ model training is completed, ultimately yielding the desired generated image.
Figure 2: The illustration of our HAIFIT architecture. Given $M$ paired sketch-image samples $\{(x_{i},y_{i})\in(X,Y)\}^{M}_{i=1}$, our goal is to map images from hand-drawn sketches domain $X$ to realistic images domain $Y$, while preserving as many line details from the sketch $X$ as possible in the generated image $\overline{Y}$.
III-B Multi-scale Feature Fusion Encoder
@@
The Multi-scale Feature Fusion Encoder (MFFE) consists of two modules: the Shallow Convolutional Module (SCM) and the Abstract Feature Representation Module (AFRM). As illustrated in Fig. 3, the SCM employs a convolutional network to learn essential global contour information, such as the structure of the object. Simultaneously, inspired by the LSTM architecture  [10], we design the AFRM to learn long-range dependencies in key areas of the input sketch from various perspectives, which can effectively capture style features and geometric features of strokes. By combining these features, we can effectively capture the consistency of objects’ complete structure and intricate textures within the latent space.
@@
As shown in Fig. 2, in the bottom layer, only the sketch features are inputted. In the other layers, we perform element-wise addition between the generated image features and the sketch features, and then input the resulting summed features into the MFFE. In the SCM we employ a sequence of convolutional layers, denoted as $Con$, to extract the global contour features and shape information $\widetilde{x}$ from the input sketch $x_{s}$. Here, $x_{s}\in\mathcal{R}^{b\times 3\times m\times m}$, $\widetilde{x}\in\mathcal{R}^{b\times 256\times m/4\times m/4}$, $b$ is the batch size, and $m$ depends on the network layer, so we can express this process as $\widetilde{x}=Con(x_{s})$. Specifically, the SCM extracts features in a blockwise manner, which will disrupt the continuity of strokes and result in pixel-level contour features. Since sketches are continuous and composed of multiple lines in different directions, it is crucial to consider the correlation between adjacent lines in order to effectively learn the line features of complex sketches.
To address this issue, we introduce the AFRM to learn the feature $\widetilde{x}$. The AFRM consists of a two-layer LSTM network, capturing abstract drawing intent features from four directions. After downsampling $\widetilde{x}$, we obtain the feature $\widetilde{x^{\prime}}\in\mathcal{R}^{b\times 256\times 4\times 4}$. We then swap feature positions after merging its second and third dimension features to obtain $\mathcal{R}^{4\times b\times 1\times 1024}$ and divide it into four parts, namely $\widetilde{x}_{0}$, $\widetilde{x}_{1}$, $\widetilde{x}_{2}$, and $\widetilde{x}_{3}$, with dimensions $\mathcal{R}^{b\times 1\times 1024}$. These features are then individually sequence inputted into the LSTM network, generating features $\overline{x}_{0}$, $\overline{x}_{1}$, $\overline{x}_{2}$, and $\overline{x}_{3}$ by recursively processing the input features from different directions. Next, we concatenate the features $\overline{x}_{0}$, $\overline{x}_{1}$, $\overline{x}_{2}$, and $\overline{x}_{3}$ to obtain the feature $\overline{x^{\prime}}\in\mathcal{R}^{4\times b\times 1\times 1024}$. Afterward, we restore the dimension of $\overline{x^{\prime}}$ to the dimension of $\widetilde{x^{\prime}}$. Finally, we apply a 1$\times$1 convolutional layer along the channel dimension to obtain the output feature $\overline{x}\in\mathcal{R}^{b\times 256\times 4\times 4}$.
@@
$\begin{split}\overline{x}_{dir}^{\prime}&=Cat(LSTM[\widetilde{x}_{0},\widetilde{x}_{1},\widetilde{x}_{2},\widetilde{x}_{3}])\\ \overline{x}_{rev}^{\prime}&=Cat(LSTM[\widetilde{x}_{0},\widetilde{x}_{1},\widetilde{x}_{2},\widetilde{x}_{3}])\\ \overline{x}&=Conv_{1\times 1}(Cat(\overline{x}_{dir}^{\prime},\overline{x}_{rev}^{\prime}))\end{split}$ (1)
@@
III-C Cross-level Skip Connection Module
@@
In order to effectively capture intricate details in the key areas of the sketch, we employ a pyramid generative learning approach [13] with a Cross-level Skip Connection Module (CSCM). Our model comprises a pyramid of generators $G_{n}\in\{G_{1},G_{2},\dots,G_{N}\}$, which are trained using a set of multi-scale images $x_{i}$ represented as ${x^{1}_{i},x^{2}_{i},\dots,x^{N}_{i}}$. Each generator $G_{n}$ is responsible for generating a realistic image $\overline{y}^{n}_{i}$. By adversarial training, the generator $G_{n}$ aims to deceive the associated discriminator $D_{n}\in\{D_{1},D_{2},\dots,D_{N}\}$, and the $D_{n}$ is trained to distinguish between the generated realistic images and real images. The generation of a realistic image sample starts at the coarsest scale $64\times 64$ and sequentially passes through all generators, progressing toward the finest scale $256\times 256$. At the coarsest scale, the MFFE takes a minibatch of the sketch feature maps $X_{n}$ as input. The effective receptive field at this level is typically half the height of the finer image. Each of the generators $G_{n}$ at finer scales adds details that are not generated by the previous scales. Therefore, in addition to the feature maps $X_{n}$, each multi-scale feature fusion encoder accepts an upsampled version of the generated image features from the coarser scale. Here, $X_{n}\in\mathcal{R}^{b\times 3\times m\times m}$, $Y_{n-1}\in\mathcal{R}^{b\times 3\times m/2\times m/2}$. This pyramid structure allows us to capture features at multiple scales and better handle the complex details of the sketch.
$Y_{n}=\begin{cases}G_{n}(X_{n}),&\mbox{if {n} = 1}\\ G_{n}(X_{n},Y_{n-1}),&\mbox{if {n} > 1}\end{cases}$ (3)
@@
The generator $G_{n}$ consists of a residual block and multiple convolutional layers. In the coarsest layer, the residual block comprises four convolutional layers. In addition, in the pyramid network, as the number of pixels increases from bottom to top, the residual block progressively adds one convolutional layer at each level. This progressive increment facilitates the learning of sketch feature representations. Through numerous experiments, it has been observed that incorporating the CSCM between different layers in the pyramid network can enhance the quality of the generated images, which is illustrated in Fig. 4. In detail, We extract the output feature ${X^{\prime}}_{n-1}$ from the last layer of the residual block in generator $G_{n-1}$ and concatenate it with the output feature ${X}_{n}$ from the $n$-th layer of the MFFE. Subsequently, we use $\overline{X}_{n}$ as the input for generator $G_{n}$, where $\overline{X}_{n}\in\mathcal{R}^{b\times 256\times m\times m}$. In principle, the multi-level features of the same synthetic image should exhibit similarity. By employing cross-level feature operations, we can extract similar details from the features generated by different level generators and incorporate them as attention maps into the new feature map $X_{n}$. This facilitates image translation, resulting in more detailed and high-quality images.
@@
Figure 4: An overview of the Cross-level Skip Connection Module (CSCM).
III-D Loss Function